OpenStudy (anonymous):

A company introduces a new product for which the number of units sold S is S(t)= 200 ( 5- 9/(2+t)) where t is the time in months (a) Find the avg. value of S(t) during the first year. (b) During what month does S'(t) equal the avg value during the first year?

OpenStudy (anonymous):

Either tricky question or question with really long solution..... My method of solving will be, substituting value of t, 1 through 12 , adding all the results and dividing by 12.........Don't know if there is shortcut

OpenStudy (anonymous):

I think you're supposed to use the mean value theorem. I mostly got it, but I don't know whether to use S(1) or s(0) in place of f(a) in the theorem.

OpenStudy (a_clan):

(a) S(12)/12

OpenStudy (anonymous):

my final answer for a was 5150 divided by 7. I found avg velocity by finding s(1) +s(12) all over 2. Is that wrong?

OpenStudy (anonymous):

thatz correct, i guess

OpenStudy (agreene):

\[\frac{\Delta s(t)}{\Delta t}\] would be how a physicist would find part A. \[\frac{s(12)-s(1)}{12}\] Of course an engineer would say: \[\frac{s(1)+s(2)+s(3)+...+s(10)+s(11)+s(12)}{12}\] and most mathematicians would say that there is not enough information the rest would use Harmonic Mean just to make everyone go wtf?

OpenStudy (anonymous):

haha I like your answer but I am none of those things, just a student. The physicist one looks the most "right" though, so I'll go with that,

OpenStudy (agreene):

The engineer way is Arithmetic Mean, and is generally what people mean when they ask a mean... hows that for a messed up sentence!

OpenStudy (anonymous):

lol(: thanks for your help.

OpenStudy (agreene):

And you are definitely right about using Mean Value Theorem... I just don't remember how to do that O.o I know it involves some stupid limit I dont like doing.